I cut a 28-ft wire into two pieces. One piece is 20ft shorter than the other, how long is each piece?
Answer by swincher4391(1107)

Call these two pieces x and y.
Then x +y = 28
x = y - 20
Use subst.
(y-20) + y = 28
2y -20 = 28
2y = 48
y = 24
So x = 4.
The two pieces are 4ft and 24ft.